# Perception-Distortion Balanced Super-Resolution: A Multi-Objective Optimization Perspective

### v1

@@ I Introduction
Image super-resolution (SR) [1] is a valuable technique to reconstruct a high-resolution (HR) image with good quality from a low-resolution (LR) input image. Benefiting from the deep learning techniques, it has become popular to train a deep neural network (DNN) for SR by using a large amount of LR-HR image pairs [2, 3, 4, 5, 6, 7, 8]. One commonly used loss function to optimize the SR models is the distortion (or fidelity) measure, e.g., the norm of the prediction errors [2, 9, 10, 11]. However, SR is a typical ill-posed problem, and there are many possible HR estimations for the given LR input. It is well-known that the $\ell_{2}$ or $\ell_{1}$ loss tends to generate over-smoothed HR estimates [12, 13] although they can result in good distortion measures such as peak-to-noise ratio (PSNR).
@@
To obtain perceptually more realistic HR images, the perception-distortion trade-off methods [19, 7, 8, 20, 21, 22] have been developed in SR. Usually, they train a generative adversarial network (GAN) [23] by using the adversarial loss. Instead of minimizing the pixel-wise $\ell_{2}$ or $\ell_{1}$ loss or local structure-wise SSIM or perceptual loss, the adversarial loss is optimized to minimize the distribution divergence between the predicted SR images and HR images. In practice, it is unstable and even impossible to train the SR model by using only the adversarial loss. Therefore, most of the GAN-based SR methods weigh the $\ell_{1}$, perceptual and adversarial losses [7, 8] into one single loss for optimization.
While GAN-based SR methods can generate a certain amount of realistic details, they will also introduce undesired artifacts [24, 17]. Many following works [25, 12, 26, 27, 28] aim to reduce the artifacts caused by the adversarial loss while keeping the realistic details. Dario et al[25] applied $\ell_{1}$ and adversarial losses in both spatial and Fourier domains to improve the restoration of high-frequency details. Liang et al[12] designed an LDL loss to suppress GAN-generated artifacts by considering their statistical difference from visually friendly details. To avoid artifacts in low-frequency areas, a region-aware adversarial loss [26] was designed to help the discriminator pay more attention to high-frequency areas. Zhang et al[27] used a similar idea to handle the low-frequency and high-frequency areas differently by using a constrained loss. Xie et al[28] proposed to detect the artifact regions and developed a finetuning procedure to improve GAN-based SR models.
@@ II-B Perception-oriented SR Methods
To make the reconstructed SR images look more perceptually realistic, SRGAN [7] combined the adversarial loss with distortion-oriented losses (e.g., $\ell_{1}$ loss) in training. Following SRGAN, ESRGAN [8] used the VGG features before activation and employed the RRDB backbone [8] as the generator model, improving much the perceptual quality. While SRGAN and ESRGAN can generate new image details, they introduce many unnatural visual artifacts because of the unstable adversarial training. Therefore, many of the following works aim to reduce the GAN-generated artifacts from the perspectives of frequency division [26, 25, 27, 44], or different statistical properties [12, 28]. In addition, some works [21] aimed to optimize the no-reference image quality metrics, such as NIQE [45] and Ma [46], which are non-differentiable. For example, RankSRGAN [21] designed a differentiable rank-content loss to rank the quality of predictions with NIQE or Ma index.
@@ II-C Evolutionary Algorithms in DNN Learning
EA [32, 33] is a classical search-based optimization technique, where individuals interactively evolve from a population. The EA operators mainly consist of crossover [47], mutation, and selection. Due to the global search and gradient-free properties, EA has proven its effectiveness in many applications [48, 49, 50]. Recently, it has also been applied to DNN learning [51, 52, 53]. Guo et al[51] applied evolutionary architecture search to construct a simplified supernet with all architectures optimized simultaneously. Yu et al[52] searched over a big single-stage model that contains small children models. EAGAN [53] applied an efficient two-stage EA-based network search framework to learn GANs. Another application is network compression. For example, Zhou et al[54] established filter pruning as a multi-objective optimization problem and proposed a knee-guided EA algorithm to trade-off between the scale of parameters and performance. In addition, EA algorithm can be used as the network optimizer. ESGD [55] alternately applied SGD and EA in one framework to improve the average performance of DNN. Zhang et al[56] proposed a hierarchical cluster-based suppression algorithm to remove similar weights and improve population diversity. GEMONN [57] determined the search direction in the EA optimizer by the gradient of weights to improve the efficiency of EA for training DNN.
@@ III The Proposed Method
III-A Multi-Objective Formulation
${L_{GAN}}={\alpha_{1}}{L_{pix}}+{\alpha_{2}}{L_{percep}}+{\alpha_{3}}{L_{adv}},$ (1)
@@ III-B The Hybrid EA-Adam Optimizer
The Adam step is illustrated in Fig. 2(b). Initialized by a pre-trained model with $\ell_{1}$ loss, the other $N-1$ generator-discriminator pairs are updated by Adam separately, as shown in the Adam step of Table I. We fix the first model across the training process as it has the best reconstruction fidelity by training with pure pixel-wise $\ell_{1}$ loss. Then, the EA step starts, as illustrated in Fig. 2(c). Note that updating the discriminator networks in EA or not has little impact on the final SR performance. This may be because the EA step focuses on the model divergence, while the adversarial training facilitates the model convergence in the Adam step. Therefore, the discriminator models are only updated in Adam steps to improve training efficiency.
@@
To select a good model from the parent models and the newly-generated offspring models, we measure the model performance by using an aggregation function derived from the Tchebycheff decomposition [60] as follows:
@@
where $z_{1}^{*}$ and $z_{2}^{*}$ are the moving minimum values of $f_{1}$ and $f_{2}$ recorded until the current iteration, and weights $\lambda$ are uniformly distributed within $(0,1]$. Specifically, we have $\lambda={0,\frac{1}{{N-1}}},{\frac{2}{{N-1}},\cdots,1}$ for the $N$ models, respectively. Note that the momentum in the Adam optimizer should be initialized after the EA optimization step. Finally, a set of generator models with different perception-distortion preferences are obtained to form a PF, as shown in Fig. 1. We will see that the SR models optimized by the hybrid EA-Adam optimizer can achieve both better perceptual quality and reconstruction fidelity than the corresponding models optimized by Adam, as validated in Section IV-C.
@@ IV-A Experiment Setup
Datasets and evaluation metrics. We use the DIV2K dataset [62] to train the population of SR models. We use 700 images in DIV2K for training the weight regression network and the remaining 100 images as the validation data in the network fusion process. We evaluate the performance of different methods on 6 benchmarks, including Set5 [63], Set14 [64], BSD100 [65], Urban100 [66], Manga109 [67], and DIV2K100 [62]. We compute the PSNR and SSIM [14] indices on the Y channel in the YCbCr space to measure the distortion degree and compute the LPIPS [30] index in the RGB space to evaluate the perceptual quality.
Backbones and compared methods: As in many previous perceptually realistic SR works [21, 20, 12, 26], we adopt SRResNet [7] and RRDB [8] as the backbone networks. SRResNet is a lightweight network proposed in SRGAN [7], while RRDB [8] is widely used in later GAN-based SR methods [20, 68, 12] for its superior performance. We compare our two models with the methods using the SRResNet and RRDB backbones, respectively. Specifically, we compare our SRResNet-based model with SRGAN [7] and RankSRGAN [21], and compare our RRDB-based model with ESRGAN [8], USRGAN [20], SPSR [68], LDL [12] and CAL-GAN [69]. For the methods that do not use these two backbones, we compare them with our model of similar capacity. Specifically, we compare our SRResNet-based model with CFSNet [70] and G-MGBP [19], and compare our RRDB-based model with EPSR[71] and PD-ADMM [27]. In addition, we compare SwinIR [40] trained with the $L_{GAN}$ loss by Adam optimizer and our proposed hybrid EA-Adam optimization method to validate the effectiveness of our method on the Transformer backbone. The results of the compared methods are obtained by using the officially released models. Note that we do not compare our method with the distortion-oriented SR methods, because this work focuses on perception-distortion balance.
@@
TABLE II: {justify} Quantitative comparison between the state-of-the-art methods and the proposed hybrid EA-Adam optimization method. Two groups of methods trained with the same backbone are compared, where SRResNet [7] and RRDB [8] backbones are used, respectively. The best results of each group are highlighted in bold.
Method
SRGAN [7]
RankSR- GAN [21]
SRResNet + Ours
ESRGAN [8]
USRGAN [20]
SPSR [68]
LDL [12]
CAL- GAN [69]
RRDB + Ours
Set5
PSNR $\uparrow$
29.93
29.77
30.25
30.45
30.85
30.40
30.99
31.04
31.31
@@
0.1254
0.1318
0.1208
0.1151
0.1284
0.1126
0.1008
0.1072
0.1038
#FOLPs 10.38G 73.43G
#Params 1.52M 16.69M
TABLE III: {justify} Quantitative comparisons between the state-of-the-art perceptually realistic SR models and the models trained by our method. The ’RRDB+Ours-PD’ is based on the RRDB backbone with the weight setting in PD-ADMM [27]. The ‘SwinIR-$L_{GAN}$’ is the SwinIR model trained with the $L_{GAN}$ loss by the Adam optimizer. The best results of each group are highlighted in bold.
Method
G-MGBP  [19]
CFSNet [70]
SRResNet +Ours
EPSR  [71]
RRDB +Ours
PD-AD MM[27]
RRDB +Ours-PD
SwinIR -$L_{GAN}$ [40]
SwinIR +Ours
Set5
@@
0.1502
0.1619
0.1330
0.1518
0.1239
0.1401
0.1336
0.1268
0.1268
BSD100
@@ IV-B Comparison with State-of-the-Arts
From Table II, we can see that for the lightweight network models with the SRResNet backbone, SRGAN gets better PSNR and SSIM scores than RankSRGAN on almost all the benchmarks, while they have higher or lower LPIPS scores on different benchmarks. Our proposed method improves both the fidelity- and perceptual-oriented metrics by a large margin under all benchmarks without increasing the SR model complexity. This validates the advantages of our hybrid EA-Adam optimization strategy, which could push the PF toward the optimal point of solution space. For models trained with the RRDB backbone, we can see that our proposed method achieves better results than the competing methods ESRGAN, USRGAN, SPSR, LDL, and CAL-GAN in most cases. The LDL also demonstrates competing results in several cases, especially for the perceptual quality metric on the Urban100 dataset. This advantage comes from its iterative and local discriminative strategy that can effectively inhibit visual artifacts and encourage the generation of high-frequency details.
From Table III, we can see that the proposed method achieves consistent improvement on most benchmarks in terms of both perceptual quality (LPIPS) and reconstruction accuracy (PSNR and SSIM), although the compared methods G-MGBP, CFSNet, and EPSR have more model parameters and FLOPs (e.g., 22.42G FLOPs and 19.22G FLOPs for ‘G-MGBP’ and ‘CFSNet’, while 10.38G FLOPs for ‘SRResNet+Ours’). In addition, we test the released model of PD-ADMM, which performs favorably in terms of distortion control yet sacrifices the perceptual quality. Since PD-ADMM assigns equal weights, i.e., [0.5, 0.5], to the two objectives, we also apply this setting in Eq. (2). One can see that our method outperforms PD-ADMM in almost all cases but with fewer FLOPs and parameters. Furthermore, our method with the SwinIR backbone also demonstrates significant improvement over its counterparts, which verifies that our method is applicable to both CNN-based and transformer-based models.
@@
We provide 4× SR visual comparisons between models (with SRResNet [7] and RRDB [8] backbones) trained by our method and other state-of-the-arts models in Fig. 4 and Fig. 5. We do not provide visual comparisons with the early published ESRGAN method due to limited space. We give the visual comparisons between PD-ADMM [27] and ‘RRDB+Ours-PD’, i.e., the RRDB model trained by our EA-Adam optimization method with the weight setting in PD-ADMM, in Fig. 6. The visual comparisons of models with the SwinIR [40] backbone are shown in Fig. 7. One can see that the proposed method can restore more correct visual patterns (e.g., the lines of zebra crossing in the third group of Fig. 5) and simultaneously inhibit the artifacts (e.g., the windows of building in the third group of Fig. 4). At the same time, our model generates many high-frequency details which facilitate perceptual quality (e.g., the lines of wood floor in the third group of Fig. 6).

### v2

@@ I Introduction
Image super-resolution (SR) [1] is a valuable technique to reconstruct a high-resolution (HR) image with good quality from a low-resolution (LR) input image. Benefiting from the deep learning techniques, it has become popular to train a deep neural network (DNN) for SR by using a large amount of LR-HR image pairs [2, 3, 4, 5, 6, 7, 8, 9]. One commonly used loss function to optimize the SR models is the distortion (or fidelity) measure, e.g., the norm of the prediction errors [2, 10, 11, 12]. However, SR is a typical ill-posed problem, and there are many possible HR estimations for the given LR input. It is well-known that the $\ell_{2}$ or $\ell_{1}$ loss tends to generate over-smoothed HR estimates [13, 14] although they can result in good distortion measures such as peak-to-noise ratio (PSNR).
@@
To obtain perceptually more realistic HR images, the perception-distortion trade-off methods [20, 7, 8, 21, 22, 23] have been developed in SR. Usually, they train a generative adversarial network (GAN) [24] by using the adversarial loss. Instead of minimizing the pixel-wise $\ell_{2}$ or $\ell_{1}$ loss or local structure-wise SSIM or perceptual loss, the adversarial loss is optimized to minimize the distribution divergence between the predicted SR images and HR images. In practice, it is unstable and impossible to train the SR model using only the adversarial loss. Therefore, most GAN-based SR methods weigh the $\ell_{1}$, perceptual and adversarial losses [7, 8] into one single loss for optimization.
While GAN-based SR methods can generate some realistic details, they will also introduce undesired artifacts [25, 18]. Many following works [26, 13, 27, 28, 29, 30, 31] aim to reduce the artifacts caused by the adversarial loss while keeping the realistic details. Dario et al[26] applied $\ell_{1}$ and adversarial losses in both spatial and Fourier domains to improve the restoration of high-frequency details. Liang et al[13] designed an LDL loss to suppress GAN-generated artifacts by considering their statistical difference from visually friendly details. To avoid artifacts in low-frequency areas, a region-aware adversarial loss [27] was designed to help the discriminator pay more attention to high-frequency areas. Zhang et al[28] used a similar idea to handle the low-frequency and high-frequency areas differently by using a constrained loss. Xie et al[29] proposed to detect the artifact regions and developed a finetuning procedure to improve GAN-based SR models.
@@ II-B Perception-oriented SR Methods
To make the reconstructed SR images perceptually more realistic, SRGAN [7] combined the adversarial loss with distortion-oriented losses (e.g., $\ell_{1}$ loss) in training. Following SRGAN, ESRGAN [8] used the VGG features before activation and employed the RRDB backbone [8] as the generator model, improving the perceptual quality. DualFormer [31] leveraged the spectral and spatial discriminators for identifying high-frequency and low-frequency information, respectively. While GAN-based methods can generate new image details, they may introduce many unnatural visual artifacts because of unstable adversarial training. Therefore, many of the following works aim to reduce the GAN-generated artifacts from the perspectives of frequency division [27, 26, 28, 47, 48], the weight of layers in perceptual loss [30, 49], or different statistical properties [13, 29]. In addition, some works [22] aim to optimize the no-reference image quality metrics, which can be non-differentiable. For example, RankSRGAN [22] designed a differentiable rank-content loss to rank the quality of predictions with NIQE [50] or Ma [51] index.
@@ II-C Evolutionary Algorithms in DNN Learning
EA [35, 36] is a classical search-based optimization technique, where individuals interactively evolve from a population. The EA operators mainly consist of crossover [52], mutation, and selection. Due to the global search and gradient-free properties, EA has proven its effectiveness in many applications [53, 54, 55]. It has recently been applied to DNN learning [56, 57, 58]. Guo et al[56] applied evolutionary architecture search to construct a simplified supernet with all architectures optimized simultaneously. Yu et al[57] searched over a big single-stage model that contains small children models. EAGAN [58] applied an efficient two-stage EA-based network search framework to learn GANs. Another application is network compression. For example, Zhou et al[59] established filter pruning as a multi-objective optimization problem and proposed a knee-guided EA algorithm to trade-off between the scale of parameters and performance. In addition, EA algorithm can be used as the network optimizer. ESGD [60] alternately applied SGD and EA in one framework to improve the average performance of DNN. Zhang et al[61] proposed a hierarchical cluster-based suppression algorithm to remove similar weights and improve population diversity. GEMONN [62] determined the search direction in the EA optimizer by the gradient of weights to improve the efficiency of EA for training DNN.
@@ III The Proposed Method
III-A Multi-Objective Formulation
${L_{GAN}}={\alpha_{1}}{L_{pix}}+{\alpha_{2}}{L_{percep}}+{\alpha_{3}}{L_{adv}},$ (1)
@@ III-A Multi-Objective Formulation
Figure 3: Illustration of the convergence process of the proposed EA-Adam optimizer in an EA-Adam cycle ($T_{Adam}+T_{EA}$). The Adam steps primarily focus on optimization convergence, enabling rapid search for improved objective losses. In contrast, the EA step emphasizes divergence optimization, ensuring a uniform spread of solutions across the plane.
@@ III-B The Hybrid EA-Adam Optimizer
The Adam step is illustrated in Fig. 2(b). Initialized by a pre-trained model with $\ell_{1}$ loss, the other $N-1$ generator-discriminator pairs are updated by Adam separately, as shown in the Adam step of Table I. We fix the first model across the training process as it has the best reconstruction fidelity by training with pure pixel-wise $\ell_{1}$ loss. Then, the EA step starts, as illustrated in Fig. 2(c). The adversarial training facilitates the model convergence in the Adam step to obtain better objective scores, while the EA step focuses on the model divergence, enabling the optimized models uniformly distributed on the performance plane spanned by the two objective indices. Therefore, the discriminators are only updated in the Adam steps to enhance convergence.
@@
To select a good model from the parent models and the newly generated offspring models, we measure the model performance by using an aggregation function derived from the Tchebycheff decomposition [65] as follows:
@@
where $z_{1}^{*}$ and $z_{2}^{*}$ are the moving minimum values of $f_{1}$ and $f_{2}$ recorded until the current iteration, and weights $\lambda$ are uniformly distributed within $(0,1]$. Specifically, we set $\lambda={0,\frac{1}{{N-1}}},{\frac{2}{{N-1}},\cdots,1}$ for the $N$ models, respectively. With different aggregation loss weights $\lambda$, the multiple optimization problem can be divided into $N$ sub-problems of single-objective optimization for different perception-distortion preferences. Note that the momentum in the Adam optimizer should be initialized after the EA optimization step. Finally, a set of generator models with different perception-distortion preferences are obtained to form a PF, as shown in Fig. 1.
To illustrate the training convergence process of the EA-Adam optimization algorithm, in Fig. 3 we demonstrate the performance of optimized models on the two conflicting objectives $f_{1}$ and $f_{2}$ over an Adam-EA cycle on the B100 dataset. Different shaped symbols represent the $N=4$ optimized models, where the green color denotes the model optimized after the first step of Adam, the blue color denotes the model optimized after $T_{Adam}$ steps by Adam ($T_{Adam}=10$ in this experiment), and the red color denotes the model further optimized by EA with $T_{EA}=1$ step. We can see that Adam focuses on the convergence of the 4 models, enabling a fast search for better objective scores of them. However, due to the opposite optimization gradient directions of $f_{1}$ and $f_{2}$, Adam suffers from optimization stagnation. That is, the optimized models targeting at different perception-distortion objectives tend to converge on the same side of the plane, making it difficult to form a diverse solution space. Fortunately, the introduction of EA enhances the diversification of the $N$ models so that the optimized models are more evenly distributed in the optimization plane with improved performance.
@@ IV-A Experiment Setup
Datasets and evaluation metrics. To train the population of SR models, we use either the DIV2K dataset [67] (800 images) or the DF2K dataset [67, 68] (800 DIV2K images and 2650 Flickr2k images) according to the training datasets of competing methods. For the network fusion process, 700 images of DIV2K are used to train the weight regression network, and the remaining 100 images are used as the validation data in the network fusion process. Note that we do not employ any test data in the stages of network fusion process. Once learned, our weight regression network is fixed and applied to any given test set.
Among the 800 images in the DIV2K training dataset, we utilize 700 images as training data for the weight regression network, and the remaining 100 images are used as the validation data in the network fusion process. We have not employed any test data during the training stage. Once learned, our weight regression network is fixed and applied to any given test set. Therefore, our method is zero-shot, which aligns with the approach taken by the referenced and compared SR methods.
Following prior arts [8, 13, 48], we conduct experiments with a scaling factor of 4×4$\times$4 × on both synthetic (downsampled using MATLAB bicubic kernel) and real-world (degraded using RealESRGAN pipeline [69]) experiments. For the bicubic-degraded SR task, we evaluate the performance of different methods on 6 benchmarks, including Set5 [70], Set14 [71], BSD100 [72], Urban100 [73], Manga109 [74], and DIV2K100 [67]. We compute the PSNR and SSIM [15] indices on the Y channel in the YCbCr space to measure the distortion degree, and compute the LPIPS [33] and DISTS [75] indices in the RGB space to evaluate the perceptual quality. For the real-world-degraded SR task, we evaluate the performance of different methods on the RealSR [76] and DrealSR [77] benchmarks.
Backbones and compared methods: As in many previous perceptually realistic SR works [22, 21, 13, 27], we adopt SRResNet [7] and RRDB [8] as the backbone networks. SRResNet is a lightweight network proposed in SRGAN [7], while RRDB [8] is widely used in later GAN-based SR methods [21, 78, 13] for its superior performance. Specifically, we compare our SRResNet-based model with SRGAN [7] and RankSRGAN [22], and compare our RRDB-based model with ESRGAN [8], SPSR [78], LDL [13], CAL-GAN [79] and WGSR [48]. In addition, Transformer-based backbones have become popular as they can capture long range dependency. Therefore. we compare SwinIR [43] trained with the $L_{GAN}$ loss by Adam optimizer and our proposed hybrid EA-Adam optimization method to validate the effectiveness of our method on the Transformer backbone. For the methods that apply different backbones or training datasets, we compare them with our model of similar capacity and training datasets. Specifically, we compare our SRResNet-based model with CFSNet [80] and G-MGBP [20], and compare our RRDB-based model with SROOE[81], PD-ADMM [28] and DualFormer [31]. We further validate the proposed method for the real-world SR tasks, and compare the obtained ‘RealESRGAN+Ours’ model with RealESRGAN [69], IKC [82], IKR-Net [83], LDL [13] and DASR [38]. The results of the compared methods are obtained using the officially released models. Note that we do not compare our method with the distortion-oriented SR methods, because this work focuses on perception-distortion balance.
@@
TABLE II: {justify} Quantitative comparison between the state-of-the-art methods and the proposed hybrid EA-Adam optimization method. Three groups of methods trained with the same backbone are compared, where SRResNet [7], RRDB [8] and SwinIR [43] backbones are used, respectively. The ‘SwinIR-$L_{GAN}$’ is the SwinIR model trained with the $L_{GAN}$ loss by the Adam optimizer. The best results of each group are highlighted in bold.
Method SRGAN [7] RankSR- GAN [22] SRResNet + Ours ESRGAN [8] SPSR [78] LDL [13] CAL- GAN [79] WGSR [48] RRDB + Ours SwinIR + $L_{GAN}$ SwinIR + Ours
Training Dataset DIV2K DIV2K DIV2K DF2K+OST DIV2K DIV2K DIV2K DIV2K DIV2K DIV2K DIV2K
Set5 PSNR $\uparrow$ 29.93 29.77 30.25 30.45 30.40 30.99 31.04 30.56 31.31 30.35 30.40
@@
#FLOPs 10.38G 73.43G 4.42G
#Params 1.52M 16.69M 0.93M
TABLE III: {justify} Quantitative comparisons between the state-of-the-art perceptually realistic SR models and the models trained by our method. The ’RRDB+Ours-PD’ employs the RRDB backbone and the weight setting in PD-ADMM [28]. The ’DualFormer+Ours’ employs the RRDB backbone and the discriminator in DualFormer [31]. The best results of each group are highlighted in bold.
Method G-MGBP  [20] CFSNet [80] SRResNet +Ours PD-AD MM[28] RRDB +Ours-PD SROOE  [30] DualFormer  [31] RRDB +Ours DualFormer +Ours
Training Dataset DIV2K DIV2K DIV2K DIV2K DIV2K DF2K DF2K DF2K DF2K
@@
Figure 9: Visual comparisons among the proposed EA-Adam optimization method with SROOE and DualFormer.
@@ IV-B Comparison with State-of-the-Arts
From Table II, we can see that for the lightweight network models with the SRResNet backbone, SRGAN gets better PSNR and SSIM scores than RankSRGAN on almost all the benchmarks, while they have higher or lower LPIPS scores on different benchmarks. Our proposed method improves the fidelity- and perceptual-oriented metrics by a large margin under all benchmarks without increasing the SR model complexity. This validates the advantages of our hybrid EA-Adam optimization strategy, which could push the PF toward the optimal point of solution space. For models trained with the RRDB backbone, we can see that our proposed method achieves better results than the competing methods ESRGAN, SPSR, LDL, CAL-GAN, and WGSR in most cases. The LDL also demonstrates competing results in several instances, especially for the perceptual quality metric on the Urban100 dataset. This advantage comes from its iterative and local discriminative strategy that can effectively inhibit visual artifacts and encourage the generation of high-frequency details. Furthermore, our method with the SwinIR backbone demonstrates significant improvement over its counterparts, which verifies that the proposed method can be applied to both CNN-based and transformer-based models.
From Table III, we can see that the proposed method achieves consistent improvement on most benchmarks in terms of both perceptual quality (LPIPS) and reconstruction accuracy (PSNR and SSIM), although the compared methods such as G-MGBP, CFSNet and SROOE have more model parameters and FLOPs (e.g., 22.42G FLOPs and 19.22G FLOPs for ‘G-MGBP’ and ‘CFSNet’, while 10.38G FLOPs for ‘SRResNet+Ours’). In addition, we test the released model of PD-ADMM, which performs favorably in terms of distortion control yet sacrifices the perceptual quality. PD-ADMM applies the same weight, i.e., [0.5, 0.5], for perception and distortion objectives. A constraint term is further introduced to ensure that the low-frequency information in two stages is equal. The ADMM optimization is applied to solve the constrained optimization problem in PD-ADMM. Since it is improper and difficult to integrate the constraint term into our proposed hybrid EA-Adam approach with $N$-optimized models, to fairly compare our method with PD-ADMM, we apply the optimization objective of PD-ADMM to EA-Adam without its constraint. One can see that our method outperforms PD-ADMM in almost all cases but with fewer FLOPs and parameters. SROOE employs an additional condition network to learn the loss aggregation weights, leading to significantly larger model size and Flops than the other methods. In contrast, our EA-Adam method is more efficient while achieving comparable results with SROOE. Our method obtains better performance than DualFormer on both perception and distortion metrics. Note that DualFormer employs a spectral and a spatial discriminator. Our EA-Adam presents a new training strategy from the optimization perspective, which can be applied to DualFormer to further enhance its performance, as shown in ‘DualFormer+Ours’ in Table III.
@@
We provide 4× SR visual comparisons between models (with SRResNet [7], RRDB [8] and SwinIR [43] backbones) trained by our method and other state-of-the-arts models in Fig. 5, Fig. 6 and Fig. 7. We do not compare with the early published method ESRGAN due to limited space. In addition, we give the visual comparisons between PD-ADMM [28] and ‘RRDB+Ours-PD’, i.e., the RRDB model trained by our EA-Adam optimization method with the weight setting in PD-ADMM, in Fig. 8. The visual comparisons between SROOE [30], DualFormer [31] and our proposed EA-Adam optimization method are presented in Fig. 9. One can see that our method can restore more correct visual patterns (e.g., the steps in the second group of Fig. 5 and the zebra crossings in the third group of Fig. 6) while inhibiting the artifacts (e.g., the building windows in the fifth group of Fig. 9). At the same time, our model can generate many high-frequency details (e.g., the lines of buildings and wood floor in the first and third group of Figs. 7 and 8), respectively.